OpenStudy (anonymous):

what is the slope -2/3 passing through point (3,5)

OpenStudy (anonymous):

What are you asking, for the line? If the slope is -2/3, and the point is (3, 5), then use this formula: (y - a) = m (x - b), where (a, b) = (3, 5) and m = -2/3